I have really been struggling with this chapter especialy section 5.6 Applications of Rational Expressions. I guess the story problems are wearing me down. could you please help me to see these in a better light? I tried setting it up however I know I am going about it wrong.
Sec 5.6 N 26
Science and medicine. A plane flies 720 mi against a steady 30-mi/h head wind and then returns to the same point with the wind. If the entire trip takes 10 h, what is the plane’s speed in still air?
(720/x+30)+ (720/x-30)=10

Answer by stanbon(48558)

A plane flies 720 mi against a steady 30-mi/h head wind and then returns to the same point with the wind. If the entire trip takes 10 h, what is the plane’s speed in still air?
-------------------------
With the wind DATA:
distance = 720 mi ; rate = p+30 mph ; time = d/r= 720/(p+30)
-----------
Against the wind DATA:
distance = 720 mi ; rate = p-30 mph ; time = d/r = 720/(p-30)
---------
EQUATION:
time + time = 10 hrs.
720/(p+30) + 720/(p-30)= 10
72/(p+30) + 72/(p-30)= 1
72p+72p = p^2-900
p^2-144p-900=0
(p-150)(p+6)=0
plane speed = 150 mph
cheers,
Stan H.